OpenStudy (anonymous):

Solve the following system and show work so I can see it plz 3y − 2x = 11 y + 2x = 9

OpenStudy (anonymous):

y+2x=9 y=9-2x 3(9-2x)-2x=11 3(9)-3(2x) 27-6x-2x=11 (combine like terms) 27-8x=11 -27 -27 -8x=-16 (divide both sides by -8) x=2 y=9-2(2) y=9-4 y=5 x=2

OpenStudy (anonymous):

Hope this is what you were looking for. Any questions?

OpenStudy (anonymous):

no thank you

OpenStudy (anonymous):

np